Week of 060911

Open Actions from last week: Move 2nd cms RB from critical area. Stop c2sc4 activity.

Chair: H.Renshall

Gmod: M.Barosso

Smod: I.Reguero

Monday:

Log: Nothing

New Actions: UPS reconfiguration tomorrow. Tim to announce the time.

Discussion: GSSDATLAS reported CASTOR failures over the weekend. Olof to investigate. They also reported large memory jobs running on small memory machines. Ulrich to follow up.

Tuesday:

Log: UPS reconfig from now for 4 hours. No physics UPS cover during this time.

New Actions: Prepare second cms gLite RB. HRR to follow up on failing GD diskserver lxfsrk524.

Discussion: cms data export is running at 150 MB/s but this is not the Tier 0 export exercise. The recent CASTOR diskserver problems are put down as due to their iptables which have been removed for the moment (was part of monitoring).

Wednesday

Log:

Actions:

Discussion:

Thursday

Log:

Actions:

Discussion:

Friday

Log:

Actions:

Discussion:

Your signature to copy/paste:

Force new revision help | | or or or Access keys: S = Save, Q = Quiet save, K = Checkpoint, P = Preview, C = Cancel

Edit | Attach | Watch | Print version | History: r4 < r3 < r2 < r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r4 - 2007-02-02 - FlaviaDonno
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LCG All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback